Solve for x.


2x^2 = -9x - 9

Answer by checkley71(8403)

2X^2+9X+9=0
(2X+3)(X+3)=0
2X+3=0
2X=-3
X=-3/2
X=-1.5 ANSWER.
X+3=0
X=-3 ANSWER.